OpenStudy (anonymous):

what is the mixed radical for squareroot 64512

zepdrix (zepdrix):

So I think a good way to approach this is to see if the number is divisible by PERFECT SQUARES. Remember what a perfect square is? 4 is a perfect square. Along with 9, 16, 25, and so on. \[\large 2^2=4, \qquad 3^2=9, \qquad 4^2=16, \qquad ...\] So I started by dividing out a 16, and it divided out evenly giving us,\[\large \sqrt{16\times4032}\]Which will simplify to,\[\large \sqrt{16}\sqrt{4032}\]\[\large 4\sqrt{4032}\]Understand the process? We can pull out a lot more than just a 16 though, so you have to try some more numbers. If you're still confused, let me know.